BIT Bank (in millions)
Assets Liabilities
Reserves $48 Deposits $340
Loans $360 Bank Capital $68

NAT Bank (in millions)
Assets Liabilities
Reserves $48 Deposits $400
Loans $360 Bank Capital $8

Assume that both BIT Bank and NAT Bank have the same net profit after tax of $8 million.

a. Calculate for each bank, BIT Bank and NAT Bank, its: i. return on assets (ROA); ii. return on equity (ROE); and iii. leverage ratio. Show all your calculations.

b. With reference to your answers in (a), which Bank (Bank BIT or Bank NAT) is more attractive for shareholders? Show your calculations to explain your answers.

c. Which bank (Bank BIT or Bank NAT) is riskier in case of loan depreciation at $60 million? Explain. Show your calculations to explain your answers.
1
Expert's answer
2020-03-25T10:32:26-0400

a)BIT Bank

"ROA=\frac{net profit}{assets}=\frac{8}{48+360}=0.019"

"ROE=\frac{net profit}{Bank Capital}=\frac{8}{68}=0.118"

"leverage ratio=\frac{deposits}{Bank Capital}=\frac{340}{68}=5"


NAT Bank:

"ROA=\frac{net profit}{assets}=\frac{8}{48+360}=0.019"

"ROE=\frac{net profit}{Bank Capital}=\frac{8}{8}=1"

"leverage ratio=\frac{deposits}{Bank Capital}=\frac{400}{8}=50"

b)Attractive to shareholders is NAT Bank.since the return on equity is higher. But this bank has a very high leverage/

с) Bank NAT is riskier in case of loan depreciation at $60 million:lower coverage

"\frac{300}{400}=0.75", Bank NAT


"\frac{300}{340}=0.88", BIT Bank